Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Bi-Weekly Remedial Activities Update
Report # 28: March 12 – March 23, 2007

« Back to Atlas Tack Superfund Site Fact Sheet

I. ACTIVITIES PERFORMED DURING THE WEEKS ENDING: 16 March & 23 March

Site Preparation
  • Improved the culvert crossing where the Solid Waste Disposal Area (SWD) haul road crosses over Boy’s Creek by installing bin blocks and bridging with 1-inch steel plate
Haul Road Construction (North of Barrier)
  • Reinforced section of SWD haul road using 4-8-inch rip rap in the section of road that was compromised from the rain even that occurred between 3/15-3/17 (3.01-inches)
  • Constructed additional haul road off of the existing SWD haul road to gain access to the marsh north of Boy’s Creek
  • Improved southern section of haul road along stockpile staging area using 1.5-3/4-inch stone to minimize soils clinging to the tires of the tractor trailers
Excavation
  • Postponed re-excavation activities south of the barrier on 3/14 due to the frost retreating in the marsh and the marsh becoming unstable to sustain heavy equipment. Focused operations north of the barrier until additional mats are mobilized to the site. Re-excavations south of the barrier will resume the week ending 3/30.
  • Continued to excavate north of the barrier east of Boy’s Creek; scheduled to be completed by 3/23
Soils Management
  • Continued to solidify only the saturated sediment on an as needed basis with saw dust
  • Continued to consolidate, stockpile, cover and secure stockpiles located in the stockpile staging area
Transportation & Disposal
  • Transported sediment to Marion landfill
  • Transported soils/sediment to Taunton landfill
  • Waiting for approval for certain stockpiles
    • Taunton – COMM-94 (±3,000 cy)
    • TREE – RCRA Subtitle D landfill (±500 cy)
    • TREE – Sub-D “Contained-In” soils (±6,000 cy)
  • Still waiting for TREE representatives to schedule a site visit
Stabilization
  • 13 stockpiles treated with the Enviroblend® mixture have been treated successfully (nine 500 cy piles and two 250 cy piles)
De-Watering
  • Continued to handle the upstream waters in Boy’s Creek by using the (3) 6-inch dry prime diesel pumps with a 3-inch trash pump for isolated locations
  • Skeleton crew comprised of (3) laborers and (1) operator worked on Saturday, 3/17 to handle the 3.01-inch storm event
  • Contacted Linda Lima (Superintendent – Fairhaven Sewer Department) on 3/20 to inform her that the site would like to begin discharge of surface water from the CID on 3/22. Ms. Lima informed WESTON that the wastewater treatment plant was operating at full capacity due to the recent storm events and that the facility could not receive additional quantities of water at this time. She suggested that they may be able to handle water the week of 3/26. In the interim, she will dispatch a representative to inspect the set-up.
Sampling
  • Collected confirmatory sediment samples from the marsh excavation located north of the hurricane barrier
  • Received analytical results from Boy’s Creek south of the barrier; creek floor has been confirmed clean, (2) sidewalls will have to be re-excavated
  • Additional air monitoring stations downwind of the stabilization effort and at the end of Egypt Lane. will continue once re-excavation activities start-up
  • Collected (10) backfill soil samples from the proposed common fill source and (2) samples from the sandy loam source located in Carver, MA
Survey
  • Continued to collect post excavation survey limits located north of the hurricane barrier
Miscellaneous
  • Continued to monitor both Adams St. and Tripp St entrances and Egypt Lane, inspecting the tractor trailers and removing clinging soils, crossing the trucks over the bike path and sweeping the roads
  • Request for approval letter from the TREE facility for “contained-in” soil acceptance requested by EPA on 3/16
  • TetraTech Rizzo personnel contracted by Atlas Tack were on site 3/9 to assess UST removal; requested copies of EPA soil data associated with the UST and that any wood debris located in the proximity of the UST be removed prior to their mobilization EPA agreed to satisfy these requests
  • Determined that sediment excavated from the marsh south of the barrier will be compared against the COMM-94 criteria
  • Evaluating potential use of soils from Marsh Island as backfill

II. POJECTED ACTIVITES FOR THE WEEKS ENDING: 30 March & 6 April

Excavation
  • Mobilize additional crane mats or dura-base mats the week of 3/26 and resume re-excavations south of the barrier
  • Begin excavation north of the barrier, west of the creek, working in a southerly direction
Backfill
  • Begin to backfill common fill and sandy loam in areas confirmed to have reached clean-up criteria in both northern and southern section of the marsh
Soils Management
  • Continue to solidify all saturated sediment excavated south and north of the hurricane barrier with saw dust
  • Continue to consolidate, stockpile, cover and secure stockpiles located in the stockpile staging area
Transportation & Disposal
  • Continue to transport COMM-94, COMM-97, COMM-97-Lined, and RCRA Subtitle D soil/sediment
Stabilization
  • Activity is complete
De-Watering
  • Continue to re-direct normal flow of water in Boy’s Creek and dewater excavation activities
  • Begin to remove standing water in the Commercial and Industrial Debris Area (CID)
Sampling
  • Continue to collect confirmatory sediment samples from excavated areas
  • Continue to collect waste characterization samples of soil/sediment stockpiles
  • Continue to collect air monitoring data
Survey
  • Continue to provide post survey of excavated areas

III. EXCAVATION SUMMARY

WEEKS/E 3/16/07 – 3/23/07
TO DATE
ACTIVITIES
INITIAL TOTAL QUANTITY (cubic yards)
INITIAL TOTAL QUANTITY
Commercial Area Excavation
Complete
2,655 CY
C.I.D. Excavation
None
7,083 CY
East Fill Excavation
None
9,046 CY
NW-Fill Area
None
8,612 CY
Lagoon Area
None
7,683 CY
South of the Barrier
285-cy
4,597-cy
North of the Barrier
3800-cy
4400-cy
Total Amount of Soils Excavated to Date:
44,076 - CY

IV. T&D BI-WEEKLY SUMMARY

DATE
DESTINATION
WASTE DETERMINATION
STOCKPILE ID
TONS
21 March 2007
Taunton
COMM-97-Lined
WF-04
±150
20 March 2007
Taunton
COMM-97-Lined
EA-25, WF-04
536.41
15 March 2007
Taunton
COMM-97-Lined
WF-016
105.39
14 March 2007
Taunton
COMM-97-Lined
MS-003, WF-016
421.64
13 March 2007
Taunton
COMM-97-Lined
MS-003
407.7
12 March 2007
Marion
COMM-97
MS-002
280.01
9 March 2007
Marion
COMM-97
MS-002, 004, 006
150
8 March 2007
Marion
COMM-97
MS-002, 004, 006
728.46

V. T&D TOTAL SUMMARY

WK/E
DESTINATION
WASTE DETERMINATION
ORIGIN
TONS
23 March 2007
Taunton
COMM-97-Lined
S. Barrier, E-Fill
± 690
16 March 2007
Taunton
COMM-97-Lined
S. Barrier, NW-Fill
934.73
16 March 2007
Marion
COMM-97
S. Barrier
280.01
9 March 2007
Marion
COMM-97
S. Barrier
878.46
9 March 2007
TREE
SUB-D
Treated soils from East Fill and soils from south of barrier
1,029.47
2 March 2007
TREE
SUB-D
Treated soils from East Fill
158.04
23 Feb 2007
Marion
SUB-D
Treated soils from East/NW Fill
672.2
16 Feb 2007
Marion
COMM-97
NW/EAST FILL
708.07
9 Feb 2007
Marion
COMM-97
LAGOON/EAST FILL
2054.98
2 Feb 2007
Marion
COMM-97
LAGOON
835.57
26 Jan 2007
Marion
COMM-97
LAGOON
1,322.25
19 Jan 2007
Marion
COMM-97
LAGOON
123.36
12 Jan 2007
Marion
COMM-97
LAGOON
1,314.80
5 Jan 2007
Marion
COMM-97
EAST FILL/NW-FILL
1,458.10
5 Jan 2007
Taunton
COMM-97-Lined
EAST FILL/NW-FILL
201.84
29 Dec 2006
Marion
COMM-97
EAST FILL/NW-FILL
972.76
29 Dec 2006
Taunton
COMM-97-Lined
EAST FILL/NW-FILL
1120.3
22 Dec 2006
Marion
COMM-97
NW-FILL
2343.75
22 Dec 2006
TREE
Sub-D Oversize
EAST FILL
168.6
15 Dec 2006
TREE
Sub-D Oversize
EAST FILL
2,006.51
15 Dec 2006
Marion
COMM-97
NW-FILL
1,134
8 Dec 2006
TREE
Sub-D Oversize
EAST FILL
1,805.81
8 Dec 2006
Marion
COMM-97
NW-FILL/CID
1,143.58
1 Dec 2006
TREE
Sub-D Oversize
EAST FILL
4043.28
24 Nov 2006
TREE
Sub-D Oversize
CID/ EAST FILL
1,590.08
17 Nov 2006
Marion
COMM-97
CID
516.55
17 Nov 2006
Bourne
COMM-97 Oversize
EAST FILL
1,964
10 Nov 2006
TREE
SUB-D
CID
1,695.56
10 Nov 2006
Bourne
COMM-97 Oversize
EAST FILL
274.67
03 Nov 2006
Marion
COMM-97
CID/EAST FILL
2,822.27
27 Oct 2006
T&D DID NOT OCCUR THIS WEEK
20 Oct 2006
T&D DID NOT OCCUR THIS WEEK
13 Oct 2006
T&D DID NOT OCCUR THIS WEEK
06 Oct 2006
Marion
COMM-97
CID
1,264.56
29 Sept 2006
Marion
COMM-97
CID
674.45
22 Sept 2006
Crapo
COMM-97
CID
436.85
15 Sept 2006
Crapo
COMM-97
CID
800.51
T&D TOTAL
39,678.27
29,825.35 CY
 

AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 8 March thru 21 March 2007 can be reviewed in the table below.

Weekly Dust Monitoring Summary
Report 28

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Church Street (µg/m3) Espinolla House Front Yard Espinolla House Rear Stairs
3/8/2007 NSC 0 0 NSC 0 0
3/9/2007 NSC 0 0 NSC 0 2
3/12/2007 NSC 3 0 NSC 0 0
3/13/2007 NSC 0 0 NSC 2 6
3/14/2007 NSC 0 0 NSC NSC NSC
3/15/2007 NSC 0 5 NSC NSC NSC
3/16/2007 NSC NSC NSC NSC NSC NSC
3/19/2007 NSC NSC NSC NSC NSC NSC
3/20/2007 NSC NSC NSC NSC NSC NSC
3/21/2007 NSC NSC NSC NSC NSC NSC

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*- Readings are for 24hr period

Jump to main content.